# thanks: Current affiliation: ISIS Neutron and Muon Source, STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, UK

### v1

@@ Abstract
A modern cold triple-axis spectrometer is planned for the High Flux Isotope Reactor (HFIR) at Oak Ridge National Laboratory. Here, we describe the conceptual principles and design of a secondary spectrometer using a multiplexed, prismatic analyzer system relying on graphite crystals and inspired by the successful implementation of the Continuous Angle Multiple Energy Analysis (CAMEA) spectrometers at the Paul Scherrer Institute [1]. This project is currently known as MANTA for Multi-analyzer Neutron Triple-Axis. We report Monte-Carlo ray-tracing simulations on a simple but realistic sample scattering kernel to further illustrate the prismatic analyzer concept’s workings, calibration, and performance.Then, we introduce a new statistical analysis approach based on the prismatic analyzer concept to improve the number of final energies measured on the spectrometer. We also study possible changes in the CAMEA design for MANTA to deliver a world-class triple-axis instrument for neutron spectroscopy of quantum condensed matter systems.
@@ I Introduction
Following the distinction made previously [1], multiplexing spectrometers can be broadly categorized into two families: wide-angle multiplexing and local multiplexing. In wide-angle multiplexing – also described as multi-channel detector systems – several analyzer channels are placed around the sample in a concentric arc, generally with fixed $\bm{k}_{f}$ ($E_{f}$). This technique increases the solid angle coverage of the analyzer system but restricts the geometry. One of the earliest examples of this concept is the MADBox spectrometer [4] at the Forschungsreaktor München (FRM), which made use of 61 in-plane analyzer channels. The FlatCone design at the Institut Laue-Langevin (ILL) [5] built upon this concept and uses 31 analyzer channels placed around the sample which scatter out of plane. A particularly efficient example of wide-angle multiplexing is the MACS spectrometer at the National Institute of Standards and Technology (NIST) [6], which uses analyzers with variable $\bm{k}_{f}$ and a combination of filters to suppress higher-order wavelengths. The local multiplexing technique places multiple analyzers on the same channel. An example is the IMPS [7] spectrometer, which places several analyzers at slightly offset angles from each other. The distinct analyzer crystals or blades direct neutrons toward a different spot on a position-sensitive detector (PSD). Similar concepts have been utilized in the SIKA, RITA-II, and PUMA spectrometers [8, 9, 10]. A particularly novel design is the HODACA concept, which uses a Rowland ”anti-focusing” effect by placing analyzers on a circumferential arc, each of which focuses towards a different detector, result a constant $E$ but large $Q$ coverage [11].
Figure 1: Ray-tracing simulation results illustrating the treatment of inelastic neutron scattering signal for a dispersing excitation of the sample. Simulations utilized $E_{i}=5.1$ meV and a spectrometer comprising eight analyzer arrays/stations per angular channel (see text for corresponding values of $E_{f}$). The simulations were treated (a) without and (b) with prismatic analysis (see Sec. IV for details on our analysis and calibration technique). The signal is projected and binned in the sample’s reciprocal space to produce a momentum-energy slice along $Q_{\parallel}$ using a perpendicular in-plane integration range of $\Delta Q_{\perp}=\pm 0.1$${}^{-1}$ around the average $Q_{\perp}=-2.0{\rm\AA}^{-1}$. $Q_{\parallel}$ and $Q_{\perp}$ refer to the momentum transfer along and orthogonal to, respectively, the viewing momentum-energy slice. The prismatic analysis technique resolves the signal by accounting for the distribution in final neutron energy for different pixels along a position-sensitive detector.
@@
Figure 2: Preliminary conceptual engineering renderings for the MANTA spectrometer at Oak Ridge National Laboratory’s High Flux Isotope Reactor. (a) Overview of the spectrometer on a newly designed NB-6 neutron guide system in the reconfigured cold neutron guide hall of HFIR. (b) Close-up view of a possible implementation of the secondary spectrometer using the CAMEA design showing four (of eight total) angular channels and eight analyzer stations per channel.
@@ II Instrument Concept and Model
The model of the primary spectrometer comprises an optimized neutron guide system with supermirror coatings that directs neutrons from the HFIR cold-neutron source (with the parameters of the existing HFIR cold-source) onto a 300 mm $\times$ 168 mm doubly focusing monochromator, see Ref. [19] for details. The sample is located 160 cm from the monochromator. An aspirational engineering rendering of the entire spectrometer in the HFIR cold guide hall is presented in Fig. 2(a). Using a focusing monochromator increases the neutron flux on the sample. However, it yields a broad divergence of the incident beam at the sample position, affecting the spectrometer’s resolution. By incorporating the guide system and the doubly focusing monochromator in our simulations (see Section III), these resolution effects can be accounted for. In some instances in Section V, we utilized a simplified version of the full primary spectrometer with an idealized, nearly monochromatic virtual source and a curved monochromator.
For the secondary spectrometer, we adapted the back-end design of CAMEA [1] to our virtual instrument, see Fig. 2(b) for an engineering diagram. The design consists of $N_{c}=8$ angular channels, each spanning $\Delta(2\theta)\!=\!7.5^{\circ}$, see Fig. 2(b). Each channel comprises $N_{s}$ analyzer stations in series, each built from $N_{b}$ blades of flat, highly-oriented pyrolytic graphite (HOPG) crystals. The analyzer blades of a given station have a mosaic of ${}^{\circ}$, equivalent to 60 arc minutes (60’), and are rotated to fulfill the Bragg condition for a given $E_{f}$. We label the analyzer station with index $s=1\cdots N_{s}$. Furthermore, analyzer stations between different channels are focused using Rowland geometry [3, 20]. Analyzer stations placed further from the sample are larger to account for decreased solid angle coverage for a given crystal size with increasing distance. Additional details of the instrument model are provided in Tab. 1. In some instances in Section V, we modify the CAMEA design by changing the number of stations $N_{s}$ and the analyzer mosaic.
@@
HOPG Segment Length (mm) 72.0 82.0 92.0 103 113 120 129 140
HOPG Segment Height (mm) 12.0 12.0 12.5 13 13.5 14 15 16
@@
Table 1: Parameters of each analyzer station in a single angular channel. Directly adapted from the CAMEA design [1, 12, 13]. The area of HOPG crystal blades of a given station increases with $E_{f}$ to account for the decrease in solid angle coverage at larger sample-analyzer distances.
@@
The analyzer blades reflect neutrons from the sample to an out-of-scattering plane direction and towards a series of horizontal position-sensitive detectors (PSDs), laid out in a radial pattern above the stations. Each detector has an active length of $\approx 0.9$ m. By focusing neutrons on specific locations on the detectors, each analyzer station generates a “strip” of neutrons associated with a particular $E_{f}$. As illustrated in Fig. 3 for the parameters of Tab. 1, this approach achieves local multiplexing by avoiding overlap between strips generated by different stations. Thus, at minima, this allows the acquisition of $N_{s}=8$ $E_{f}$ values per channel.
@@
Figure 3: Neutron ray-tracing simulations of a single channel with $N_{s}=8$ stations of $N_{b}=5$ for the parameters of Tab. 1 to illustrate the local multiplexing effect. A collimated “white” beam of neutrons with uniform spectral density between $E_{i}=1.0$ meV and $7.0$ meV is directed at the secondary spectrometer. For illustration, instead of realistic PSD tubes used later in this work, neutron counts are collected here using two-dimensional PSD detectors above the analyzer stations .
@@
Though this method of discrete angular and analyser channels to form a multiplexing instrument as a viable approach for decades, one of the advantages and innovations of the CAMEA concept is the utilization of the prismatic analyzer concept [21, 1, 12]. The foundations of prismatic analysis come from a first-order expansion of Bragg’s law from an analyzer crystal (with lattice spacing $d$). This yields a linear relationship between small deviations in incident angles $\Delta\theta$ and small deviations in reflected neutron wavelength $\Delta\lambda$ around a nominal Bragg reflection ($\sin\theta_{0}=\lambda_{0}/2d$), as illustrated in Fig. 4(a). As a result of analyzer mosaic, neutrons with proximate wavelengths from the nominal wavelength $\lambda_{0}$ are Bragg-scattered in different directions, Fig. 4(b). Using a position-sensitive detector, this effect can be advantageously used to encode small wavelength deviations into detector positions, Fig. 4(c). Since its inception, prismatic analysis has been featured in the design of the planned time-of-flight instruments BIFROST [17] and FARO [22].
@@
Figure 4: (a) Relation between small angular and wavelength deviations for a perfect analyzer blade. (b) In prismatic analysis, neutrons with proximate wavelengths/energies experience different angular deviations from fulfilling the Bragg scattering condition. (c) Using a position-sensitive detector far above the analyzer stations, different wavelengths become discernible on the detector. Note that in this diagram the 3.29 meV neutrons are scattered by two analyzers, an effect also captured in our ray-tracing simulations.
@@ III Simulations: Methods and Sample Kernels
To assess the performance and validate the concept of our spectrometer model, we employ neutron ray-tracing simulation using the McStas package [23]. McStas is a state-of-the-art tool used in designing modern neutron scattering instruments [17, 22, 16, 7, 6, 15, 1, 9, 24, 25, 26, 27, 28]. The program uses Monte Carlo techniques to simulate the behavior of neutron scattering instrument components with high accuracy. It also allows testing the instrument’s performance on various simple sample scattering kernels to mimic actual experiments. In this work, we employed McStas versions 2.7 and 3.1.
The primary spectrometer, from the cold source to 15 cm before the sample, was simulated in McStas for various choices of incident energy $E_{i}$, using the design of Ref. [19]. These results were pre-compiled to form a set of simulated incoming beam profiles for the secondary spectrometer. The latter was simulated by adapting the McStas code of the PSI-CAMEA back-end, which was generously provided to us by its designers [1]. The virtual instrument comprises $N_{b}=5$ analyzer blades per station, $N_{s}=8$ stations per channel, and $N_{c}=8$ channels (for a total continuous scattering angle coverage of $8\times 7.5^{\circ}\!=\!60^{\circ}$). While the actual spectrometer uses filters and a radial collimator after the sample, we did not include these components in our simulations to maintain a continuous coverage; including collimators in our simulations is left for future work. A 10 $\times$ 10 cm rectangular slit simulates the opening in front of the analyzer stations, allowing full illumination of the analyzers and unrestricted angular coverage. Unless otherwise noted, we used an infinitely thin HOPG analyzer with a mosaic of 60’. Note that modeling the analyzer in this manner does not take into account the background signal that arises from thermal diffuse scattering and phonons in HOPG, nor the effect of the decrease in peak reflectivity as a function of crystal mosaic [14, 29, 3]
We employed two types of sample kernels. The first virtual sample was a full cylinder of polycrystalline vanadium with a 2 cm height and a 1.5 cm radius. We used this incoherent scattering kernel to obtain calibration runs (see Section IV) of the analyzer system and the position-sensitive detectors. Then, a virtual single crystal with a similar cylindrical geometry was prepared using the McStas phonon kernel component to mimic a simple bosonic excitation (acoustic magnon or phonon) dispersing in three dimensions with a bandwidth below 2 meV and uniform intensity across the Brillouin zone. The parameters of this inelastic scattering kernel are found in Tab. 2.
@@
Simulations were conducted with the single crystal kernel to model a realistic scattering experiment covering energy transfers between $E=0$ meV and $E=2$ meV. As the detector system has an overall $2\theta$ coverage of $60^{\circ}$, the single crystal sample was measured for several positions of the detector system. Using the lowest angle of the detector systems as a reference, we measured with $2\theta_{i}$ values around $6^{\circ}$ (3 positions to account for the gap between the detectors), $66^{\circ}$ (5 positions), and $116^{\circ}$ (4 positions), for a net total of 12 positions. For each detector position, the single crystal sample was rotated in steps of $\Delta\phi=2.5^{\circ}$ and we ran the Monte-Carlo process with $>10^{8}$ neutrons for each sample orientation. This was repeated for two $E_{i}$ values, 5.1 meV and 5.27 meV, to produce the simulated results. The kernel was designed to measure the entire dispersion curve with $E_{i}=5.1$ meV, and the second incident energy of $E_{i}=5.27$ meV was selected to fill in the gaps in energy space. These results are shown in Section V.

### v2

@@ Abstract
A modern cold triple-axis spectrometer to study quantum condensed matter systems is planned for the High Flux Isotope Reactor (HFIR) at Oak Ridge National Laboratory. Here, we describe the conceptual principles and design of a secondary spectrometer using a multiplexed, prismatic analyzer system relying on graphite crystals and inspired by the successful implementation of the Continuous Angle Multiple Energy Analysis (CAMEA) spectrometers at the Paul Scherrer Institute. This project is currently known as MANTA for Multi-Analyzer Neutron Triple-Axis. We report Monte-Carlo ray-tracing simulations on a simple but realistic sample scattering kernel to further illustrate the prismatic analyzer concept’s workings, calibration, and performance. Then, we introduce a new statistical analysis approach based on the prismatic analyzer concept to improve the number of final energies measured on the spectrometer. We also study possible evolutions in the CAMEA design relevant for MANTA.
@@ I Introduction
Following the distinction made previously [1], multiplexing spectrometers can be broadly categorized into two families: wide-angle multiplexing and local multiplexing. In wide-angle multiplexing – also described as multi-channel detector systems – several analyzer channels are placed around the sample in a concentric arc, generally with fixed $\bm{k}_{f}$ ($E_{f}$). This technique increases the solid angle coverage of the analyzer system but restricts the geometry. One of the earliest examples of this concept is the MADBox spectrometer [4] at the Forschungsreaktor München (FRM), which made use of 61 in-plane analyzer channels. The FlatCone design at the Institut Laue-Langevin (ILL) [5] built upon this concept and uses 31 analyzer channels placed around the sample which scatter out of plane. A particularly efficient example of wide-angle multiplexing is the MACS spectrometer at the National Institute of Standards and Technology (NIST) [6], which uses analyzers with variable $\bm{k}_{f}$ and a combination of filters to suppress higher-order wavelengths. The local multiplexing technique places multiple analyzers on the same channel. An example is the IMPS [7] spectrometer, which places several analyzers at slightly offset angles from each other. The distinct analyzer crystals or blades direct neutrons toward a different spot on a position-sensitive detector (PSD). Similar concepts have been utilized in the SIKA, RITA-II, and PUMA spectrometers [8, 9, 10]. A particularly novel design is the HODACA concept, which uses a Rowland “anti-focusing” effect by placing analyzers on a circumferential arc, each of which focuses towards a different detector, resulting in a constant $E_{f}$ but large $Q$ coverage [11].
Figure 1: Ray-tracing simulation results illustrating the treatment of inelastic neutron scattering signal for a dispersing excitation of the sample. Simulations utilized $E_{i}=5.1$ meV and a spectrometer comprising eight analyzer arrays/stations per angular channel (see text for corresponding values of $E_{f}$). The simulations were treated (a) without and (b) with prismatic analysis (see Sec. IV for details on our analysis and calibration technique). The signal is projected and binned in the sample’s reciprocal space to produce a momentum-energy slice along $Q_{\parallel}$ using a perpendicular in-plane integration range of $\Delta Q_{\perp}=\pm 0.1$-1 around the average $Q_{\perp}=-2.0{\rm\AA}^{-1}$. $Q_{\parallel}$ and $Q_{\perp}$ refer to the momentum transfer along and orthogonal to, respectively, the viewing momentum-energy slice. The prismatic analysis technique resolves the signal by accounting for the distribution in final neutron energy for different pixels along a position-sensitive detector.
@@
Figure 2: Preliminary conceptual engineering renderings for the MANTA spectrometer at Oak Ridge National Laboratory’s High Flux Isotope Reactor. (a) Overview of the spectrometer on a newly designed NB-6 neutron guide system in the reconfigured cold neutron guide hall of HFIR. (b) Close-up view of a possible implementation of the secondary spectrometer using the CAMEA [1, 13] design showing $c=1-$ through $c=4$ (of eight total) angular channels, $N_{s}=8$ analyzer stations per channel, and $N_{b}=5$ HOPG blades per station.
@@ II Instrument Concept and Model
The model of the primary spectrometer comprises an optimized neutron guide system with supermirror coatings that directs neutrons from the HFIR cold-neutron source (with the parameters of the existing HFIR cold-source) onto a 300 mm $\times$ 168 mm doubly focusing monochromator. The primary was optimized for incident energies of 2.4 to 20 meV, see Ref. [19] for details. The sample is located 160 cm from the monochromator. An aspirational engineering rendering of the spectrometer in the HFIR cold guide hall is presented in Fig. 2(a). Using a focusing monochromator increases the neutron flux on the sample. However, it yields a broad divergence of the incident beam at the sample position, affecting the spectrometer’s resolution. By incorporating the guide system and the doubly focusing monochromator in our simulations (see Section III), these resolution effects can be accounted for. In some instances in Section V, we utilized a simplified, “toy model” design of the full primary spectrometer. The “toy model” was designed to ease computational intensity and used an idealized virtual source and curved monochromator that produced a similar spot size and energy resolution at the sample location.
For the secondary spectrometer, we adapted the back-end design of CAMEA [1] to our virtual instrument, see Fig. 2(b) for an engineering diagram. As the planned use of a beryllium filter in the scattered beam requires analyzed energies to below 5.2 meV, and geometrical constraints require the analyzed energy to be above 2.4 meV, the CAMEA design is particularly well suited for our purposes. The design consists of $N_{c}=8$ angular channels, each of which have $N_{d}=13$ detectors that span $\Delta(2\theta)\!=\!7.5^{\circ}$ per channel, see Fig. 2(b). Each channel comprises $N_{s}$ analyzer stations in series, each built from $N_{b}$ blades of flat, highly-oriented pyrolytic graphite (HOPG) crystals. The analyzer blades of a given station have a mosaic of 1, equivalent to 60 arc minutes (60’), and are rotated to fulfill the Bragg condition for a given $E_{f}$. We label the analyzer station with index $s=1\cdots N_{s}$. Furthermore, analyzer stations between different channels are focused using Rowland geometry [3, 20]. Analyzer stations placed further from the sample are larger to account for decreased solid angle coverage for a given crystal size with increasing distance. Additional details of the instrument model are provided in Tab. 1. In some instances in Section V, we modify the CAMEA design by changing the number of stations $N_{s}$ and the analyzer mosaic.
@@
HOPG Segment Length (mm) 72 82 92 103 113 120 129 140
HOPG Segment Height (mm) 12.0 12.0 12.5 13.0 13.5 14.0 15.0 16.0
@@
Table 1: The parameters of each analyzer station in a single angular channel as adapted from the CAMEA design [1, 12, 13]. The area of HOPG crystal blades of a given station increases with $E_{f}$ to account for the decrease in solid angle coverage at larger sample-analyzer distances.
@@
The analyzer blades reflect neutrons from the sample to an out-of-scattering plane direction and towards a series of horizontal position-sensitive detectors (PSDs), laid out in a radial pattern above the stations. Furthermore, the radial pattern has two vertical rows to minimize gaps between detectors. Each detector has an active length of $\approx 0.9$ m and a diameter of 12.6mm. By focusing neutrons on specific locations on the detectors, each analyzer station generates a “strip” of neutrons associated with a particular $E_{f}$. As illustrated in Fig. 3 for the parameters of Tab. 1, this approach achieves local multiplexing by avoiding overlap between strips generated by different stations. Thus, at minima, this allows the acquisition of $N_{s}=8$ $E_{f}$ values for each channel.
@@
Figure 3: Neutron ray-tracing simulations of a single channel with $N_{s}=8$ stations of $N_{b}=5$ for the parameters of Tab. 1 to illustrate the local multiplexing effect. A collimated “white” beam of neutrons with uniform spectral density between $E_{i}=1.0$ meV and $7.0$ meV is directed at the secondary spectrometer. For illustration, instead of realistic PSD tubes used later in this work, neutron counts are collected here using a rectangular PSD detector above the analyzer stations. The lower rack (cyan) and upper rack (red) of the detectors are outlined to show their coverage.
@@
Though this method of discrete angular and analyzer channels to form a multiplexing instrument has been accepted as a viable approach for decades, one of the advantages and innovations of the CAMEA concept is the utilization of the prismatic analyzer concept [21, 1, 12]. The foundations of prismatic analysis come from a first-order expansion of Bragg’s law from a finite mosaic analyzer crystal (with lattice spacing $d$). This yields a linear relationship between small deviations in incident angles $\Delta\theta$ and small deviations in reflected neutron wavelength $\Delta\lambda$ around a nominal Bragg reflection ($\sin\theta_{0}=\lambda_{0}/2d$), as illustrated in Fig. 4(a). As a result of analyzer mosaic, neutrons with proximate wavelengths from the nominal wavelength $\lambda_{0}$ are Bragg-scattered in different directions, Fig. 4(b). Using a position-sensitive detector, this effect can be advantageously used to encode small wavelength deviations into detector positions, Fig. 4(c). Since its inception, prismatic analysis has been featured in the design of the planned time-of-flight instruments BIFROST [17] and FARO [22].
@@
Figure 4: ((a) Relation between small angular and wavelength deviations for a perfect analyzer blade. (b) In prismatic analysis, neutrons with proximate wavelengths/energies experience different angular deviations from fulfilling the Bragg scattering condition. (c) Using a position-sensitive detector far above the analyzer stations, different wavelengths become discernible on the detector. Note that in this diagram the 3.29 meV neutrons are scattered by two analyzers, an effect also captured in our ray-tracing simulations.
@@ III Simulations: Methods and Sample Kernels
To assess the performance and validate the concept of our spectrometer model, we employ neutron ray-tracing simulations using the McStas package [23]. McStas is a state-of-the-art tool used in designing modern neutron scattering instruments [17, 22, 16, 7, 6, 15, 1, 9, 24, 25, 26, 27, 28]. The program uses Monte Carlo techniques to simulate the behavior of neutron scattering instrument components with high accuracy. It also allows testing the instrument’s performance on various simple sample scattering kernels to mimic actual experiments. In this work, we employed McStas versions 2.7, 3.1, and 3.4.
The primary spectrometer, from the cold source to 15 cm before the sample, was simulated in McStas for various choices of incident energy $E_{i}$, using the design of Ref. [19]. These results were pre-compiled to form a set of simulated incoming beam profiles for the secondary spectrometer. The latter was simulated by adapting the McStas code of the PSI-CAMEA back-end, which was generously provided to us by its designers [1]. The virtual instrument comprises $N_{b}=5$ analyzer blades per station, $N_{s}=8$ stations per channel, $N_{c}=8$ channels with $N_{d}=13$ detectors per channel. This results in a total continuous scattering angle coverage of $8\times 7.5^{\circ}\!=\!60^{\circ}$ and a total of $N_{s}*N_{c}*N_{d}=832$ points in $S(\vec{Q},\omega)$ measured simultaneously. While the actual spectrometer uses filters and a radial collimator after the sample, we did not include these components in our simulations to maintain a continuous coverage; including collimators in our simulations is left for future work. A 10 $\times$ 10 cm rectangular slit simulates the opening in front of the analyzer stations, allowing full illumination of the analyzers and unrestricted angular coverage. Unless otherwise noted, we used an infinitely thin HOPG analyzer with a mosaic of 60’. Note that modeling the analyzer in this manner does not take into account the background signal that arises from thermal diffuse scattering and phonons in HOPG, nor the effect of the decrease in peak reflectivity as a function of crystal mosaic [14, 29, 3].
We employed three types of sample kernels. The first virtual sample was a full cylinder of polycrystalline vanadium with a 2 cm height and a 1.0 cm radius. We used this incoherent scattering kernel to obtain calibration runs (see Section IV) of the analyzer system and the position-sensitive detectors. Then, a virtual single crystal with a similar cylindrical geometry was prepared using the McStas phonon kernel component to mimic a simple bosonic excitation (acoustic magnon or phonon) dispersing in three dimensions with a bandwidth below 2 meV and uniform intensity across the Brillouin zone. The parameters of this inelastic scattering kernel are found in Tab. 2. Finally, we made use of a single crystal elastic Bragg scattering sample with a mosaic of 10’, lattice constant c = 6.71 $\,\mathrm{\AA}$, and the same geometry as the inelastic scattering kernel, which produced a single Bragg peak at $\vec{Q}=(0,0,2)$ ($|\vec{Q}|=1.85\,\mathrm{\AA^{-1}}$). This sample was used to calculate the elastic $\vec{Q}$ broadening due to the instrument.
@@
The simulations with the inelastic scattering crystal kernel were conducted to model a realistic scattering experiment covering energy transfers between $E=0$ meV and $E=2$ meV. As the detector system has an overall $2\theta$ coverage of $60^{\circ}$, the single crystal sample was simulated for several positions of the detector system. Using the lowest angle of the detector systems as a reference, we aimed to measure 5 $2\theta_{i}$ values centered around three reference points, $6^{\circ}$, $66^{\circ}$, and $116^{\circ}$. However to minimize scattering signal from non-interacting neutrons, simulations where the lowest detector angle was below $2.5^{\circ}$ or the largest detector angle was at or above $180^{\circ}$ were not simulated, leading to a net total of 12 positions. For each detector position, the single crystal sample was rotated in steps of $\Delta\phi=2.5^{\circ}$ and we ran the Monte-Carlo process with $>10^{8}$ neutrons for each sample orientation. This was repeated for two $E_{i}$ values, 5.1 meV and 5.27 meV, to produce the simulated results. The kernel was designed to measure the entire dispersion curve with $E_{i}=5.1$ meV, and the second incident energy of $E_{i}=5.27$ meV was selected to fill in the gaps in energy space. These results are shown in Section V.
Additional simulations with an elastic Bragg scattering sample were conducted to determine the elastic $\vec{Q}$ broadening as a function of $E_{i}$. To do this, the full primary and secondary spectrometer were used in a similar setup as the inelastic sample simulations. However, only one angular channel was used and a single slit with an opening of $1^{\circ}$ was placed in half the simulations. Using the known position of the Bragg peak as a reference, $2\theta$ was varied from a range of $\Delta 2\theta=\pm 12^{\circ}$ from the Bragg peak in steps of $1^{\circ}$ and sample was rotated in a range of $\Delta\phi=\pm 7$ in steps of $0.5^{\circ}$. This was repeated for three $E_{i}$ values, 3.21 meV, 4.05 meV, and 5.01 meV. The results are tabulated in Section V.